Jim's aquarium holds 3.4 cubic meters of water. Its height and width are 1.5 and meters and 0.8 meter, respectively. what is the lenth of the aquarium? round your answer to the nearedt 10th

whats the varible and write a equation. then solve and answer the question. show your work

Answer by colliefan(242)

Since volume = length * width * height
3.4 = l * 1.5 * 0.8
3.4 = l*1.2
3.4/1.2=l
2.8 meters = l or length